# Foundation Models in Medical Imaging - A Review and Outlook

### v3

@@ 5.1.4 Application of SAM for Ophthalmology
Finally, some works focus on enhancing SAM for retinal image analysis. [171] propose SAM-U, a method that enhances the reliability of SAM in medical imaging through multi-box prompts and uncertainty estimation. It uses different prompts as a form of test-time augmentation to estimate the distribution of SAM predictions. Demonstrating their methods on fundus photography images, the authors show SAM’s improved performance while also providing pixel-level uncertainty, making it particularly useful for medical image segmentation. Another work that builds on SAM is SAM-OCTA [172], a method for local segmentation in Optical Coherence Tomography Angiography (OCTA) images. It fine-tunes a pre-trained SAM model using LoRA and utilizes prompt points for local segmentation of retinal vessels, arteries, and veins. It has shown state-of-the-art performance in common OCTA segmentation tasks, particularly in local vessel segmentation and artery-vein segmentation, which were not well-solved in previous works.
@@ 5.1.5 Concluding Remarks
Compared to trends in pathology and radiology, there is a predominance of vision models over vision-language models in ophthalmology. While VLMs like FLAIR and DERETFound incorporate expert knowledge, there are no foundation models that extensively utilize report data, as seen in radiology. This distinction likely stems from differences in clinical workflows and the role of textual data in decision-making. Radiological images typically capture large anatomical structures, meaning lesions, fractures, or other abnormalities must be identified within a broader spatial context. Reports play an important role in contextualizing these findings, and incorporate for example spatial relationships, comparisons with prior scans, and clinical history [173]. In contrast, ophthalmology diagnoses are more driven by direct image interpretation, with less reliance on paired textual descriptions [174, 175]. As a result, ophthalmology has fewer large-scale image-text datasets, limiting the development of vision-language foundation models in this domain.
@@ 6 Challenges and Future Directions
Foundation models have shown remarkable promise in medical image analysis, with improved generalization and robustness compared to traditional deep learning methods [176, 177]. However, several challenges need to be addressed to ensure their safe and effective deployment in clinical settings. These challenges fall into two broad categories: (1) technical barriers in model development, and (2) practical challenges in clinical integration. In the sections below, we outline each of these areas and offer directions for future research.
@@ 6.1.1 Lack of Open-Source Large-Scale Clinical Datasets
Unlike natural image datasets, medical imaging datasets often suffer from restricted access due to privacy regulations and/or institutional policies. As seen in this Tables 1-7, the availability of datasets varies quite a lot. In fact, many of the larger datasets fall under one of the following trends: 1. large datasets that were collected from public sources, such as Twitter or PubMed, and are therefore also made open-source, or 2. large datasets that are proprietary, such as those of Virchow, Phikon, GigaPath or VisionMAE. In addition, datasets collected from the internet typically do not include full volumes for 3D imaging domains, nor different magnification levels as would be present in WSIs. This means that there still exists a lack of large, clinical datasets that are publicly available, which can limit reproducibility and collaboration. Research into directions such as federated learning and differential privacy can be beneficial for solving these challenges [178, 179].
@@ 6.1.2 Increased Computational Demands of Volumetric Data
For many of the works in this review, it could be observed that larger datasets improved model performance, and that there is a trend towards scaling up. However, scaling is particularly challenging for computationally demanding imaging modalities, such as high-resolution WSIs, and volumetric CT or MRI scans, even outside the medical imaging domain [180]. As foundation models scale up in both data and architecture, ongoing research should prioritize developing more efficient solutions specifically tailored for processing volumetric data, as seen for example in [181]
@@ 6.1.3 Pathology FMs Show Plateau in Scaling-Laws
While the challenges in scaling up data and computational resources are still ongoing, pathology FMs have seen many recent successes by scaling up, downstream task performance generally improving with increased data and model size. However, some recent studies show a performance plateau when scaling along these axes [60, 70, 182, 183, 71, 72], which can be attributed to several factors.
@@
First, simply increasing data size may not be enough; better curation is needed to create a more diverse dataset that covers a range of organs, pathologies, and staining types. For example, [75] demonstrated that Virchow’s performance could be matched with a curated dataset containing ten times fewer slides. Furthermore, Atlas [79] could even outperform Virchow despite having a much smaller pre-training dataset size, by focusing on including a much broader range of scanners, stainings and magnifications. Similarly, [74] demonstrated that a smaller ViT-B/16 model pre-trained on 4 million patches using iBOT outperformed Virchow2, a model pre-trained on 350 times more histology images, across three external validation cohorts. This underscores the importance of targeted training strategies over brute-force scaling.
@@
Furthermore, downstream task comparisons often rely on public challenge datasets, but a broader range of benchmark tasks is needed to fully assess pathology FMs. [73] found performance on public tile-level benchmarks plateauing, and [183] noted more variability in performance when using challenging biomarker prediction tasks as benchmarks. To address these challenges, frameworks such as Patho-Bench [184] and eva [185] provide standardized evaluation tools, enabling systematic benchmarking across diverse datasets, tasks, and evaluation protocols.
@@
Finally, current SSL frameworks are optimized for natural images, not pathology images, which have distinct characteristics such as lack of canonical orientation and color variation, as well as interpretation differences depending on the field of view [66]. SSL that use for example data augmentations or pretext tasks, such as colorization or rotation prediction frameworks, may need domain-specific adaptations for medical imaging data. [74] further showed that, while scaling general-purpose FMs like GigaPath, H-Optimus-0, Phikon-v2, UNI, and Virchow2 often benefits downstream tasks, these models do not excel uniformly across all organs or tasks.
Although studies have explored domain-specific adjustments (e.g., weight initialization, data augmentations, magnification levels), both in pathology [63, 186, 66, 187, 182, 188, 189, 190, 191, 192, 73] and other medical imaging domains [147, 148, 149, 140], many open questions remain about developing successful SSL algorithms for medical imaging. As foundation models for all medical imaging domains are increasingly being developed, these questions become relevant for ensuring scalability, generalizability, and the effective utilization of resources across all medical imaging domains.
@@ 6.2.1 Explainability
Foundation models for medical image analysis will need to provide transparent and interpretable results, enabling clinicians and researchers to understand how the model is making predictions and what features are driving those predictions. However, the complexity of foundation models makes it difficult to understand how they make predictions, identify biases, and explain errors, which can have serious consequences in healthcare [193]. Explainable AI techniques can help address this issue. For example, [194] introduced an interpretable decomposition approach for tumor classification, while [195] developed a controllable and interpretable system for chest X-ray diagnoses. [171] add pixel-wise uncertainty estimation to SAM, making it possible to evaluate the the reliability of segmented lesions.
@@ 6.2.2 Robustness and Domain Generalization
Foundation models in medical imaging must remain robust to variations in data acquisition, staining, scanning equipment, and institutional protocols. Without this robustness, models risk producing unreliable or misleading results in real-world deployment. For example, [196] found that many current pathology FMs encode confounding factors such as medical center-specific artifacts more strongly than relevant biological signals. They proposed the robustness index, a quantitative metric to evaluate how well model representations are driven by biological rather than spurious features.
Similarly, [197] demonstrated that a distilled pathology FM (H0-mini) outperformed larger models on robustness metrics, showing resilience to staining and scanning variability on the PLISM dataset. This highlights that model size does not always correlate with robustness, and that distillation or other targeted training techniques may help improve domain generalization.
@@ 6.2.3 Fairness, Bias and Robustness
Bias presents a significant challenge for FMs, particularly in clinical contexts where the accuracy and fairness of predictions are critical. Models trained on biased datasets risk producing biased predictions, which can disproportionately affect underrepresented populations [198]. Addressing this requires diverse and representative training data, alongside robust fairness evaluation frameworks. Studies such as [199] and [200] highlight the importance of identifying and mitigating biases in FMs. Specific approaches that aim to reduce bias include cross-lingual pre-training, as proposed by [201], which helps reduce bias in vision-language models. These solutions are vital to ensure FMs provide equitable benefits across diverse patient populations.
Furthermore, ensuring the robustness of foundation models to variables like scanner type, staining protocols, and institutional differences is crucial for their generalizability across clinical settings. [196] found that many current pathology foundation models encode confounding factors, such as medical center-specific staining artifacts, more strongly than relevant biological signals like tissue or cancer type. To address this, they proposed the robustness index, a quantitative metric that measures the extent to which model representations are driven by biological features rather than spurious confounders, providing a practical tool for evaluating model reliability. [197] explore the distillation of a large foundation model into the much smaller H0-mini. Their model, demonstrates excellent robustness to variations in staining and scanning conditions, significantly outperforming other current pathology FMs.
Finally, with the growing trend of scaling up foundation models, there is a risk of making them inaccessible to institutions with limited computational resources. It should therefore be encouraged to develop more efficient FM architectures to improve accessibility [162, 67]. Additionally, parameter-efficient fine-tuning strategies, as highlighted by [202], offer a promising pathway for enabling widespread adoption while minimizing computational demands.
@@ 6.2.4 Regulation
For FMs to be successfully adopted in clinical practice, their risks must be addressed through comprehensive regulation. Medical imaging FMs, like other AI systems, can produce unreliable outputs, such as hallucinations; plausible but incorrect predictions that could lead to misdiagnoses or inappropriate treatments. Recent studies propose methods for detecting hallucinations in medical vision-language FMs [203, 204], but more research is needed to extend these approaches to vision-based FMs.
Continual learning is a promising research direction to ensure that FMs can adapt to new tasks, disease variants, or data without forgetting previously learned knowledge [205, 206]. This adaptability is crucial in the rapidly evolving medical field.
7 Statements and Declarations
All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.

### v4

@@ 5.1.4 Application of SAM for Ophthalmology
Finally, some works focus on enhancing SAM for retinal image analysis. [deng_sam-u_2023] propose SAM-U, a method that enhances the reliability of SAM in medical imaging through multi-box prompts and uncertainty estimation. It uses different prompts as a form of test-time augmentation to estimate the distribution of SAM predictions. Demonstrating their methods on fundus photography images, the authors show SAM’s improved performance while also providing pixel-level uncertainty, making it particularly useful for medical image segmentation. Another work that builds on SAM is SAM-OCTA [wang_sam-octa_2023], a method for local segmentation in Optical Coherence Tomography Angiography (OCTA) images. It fine-tunes a pre-trained SAM model using LoRA and utilizes prompt points for local segmentation of retinal vessels, arteries, and veins. It has shown state-of-the-art performance in common OCTA segmentation tasks, particularly in local vessel segmentation and artery-vein segmentation, which were not well-solved in previous works.
@@ 5.1.5 Concluding Remarks
Compared to trends in pathology and radiology, there is a predominance of vision models over vision-language models in ophthalmology. While VLMs like FLAIR and DERETFound incorporate expert knowledge, there are no foundation models that extensively utilize report data, as seen in radiology. This distinction likely stems from differences in clinical workflows and the role of textual data in decision-making. Radiological images typically capture large anatomical structures, meaning lesions, fractures, or other abnormalities must be identified within a broader spatial context. Reports play an important role in contextualizing these findings, and incorporate for example spatial relationships, comparisons with prior scans, and clinical history [pesapane2023advancements]. In contrast, ophthalmology diagnoses are more driven by direct image interpretation, with less reliance on paired textual descriptions [lee2021recommendations, shweikh2023growing]. As a result, ophthalmology has fewer large-scale image-text datasets, limiting the development of vision-language foundation models in this domain.
@@ 6 Challenges and Future Directions
Foundation models have shown remarkable promise in medical image analysis, with improved generalization and robustness compared to traditional deep learning methods [nguyen2023distribution, azizi_robust_2023]. However, several challenges need to be addressed to ensure their safe and effective deployment in clinical settings. These challenges fall into two broad categories: (1) technical barriers in model development, and (2) practical challenges in clinical integration. In the sections below, we outline each of these areas and offer directions for future research.
@@ 6.1.1 Lack of Open-Source Large-Scale Clinical Datasets
Unlike natural image datasets, medical imaging datasets often suffer from restricted access due to privacy regulations and/or institutional policies. As seen in this Tables 1-7, the availability of datasets varies quite a lot. In fact, many of the larger datasets fall under one of the following trends: 1. large datasets that were collected from public sources, such as Twitter or PubMed, and are therefore also made open-source, or 2. large datasets that are proprietary, such as those of Virchow, Phikon, GigaPath or VisionMAE. In addition, datasets collected from the internet typically do not include full volumes for 3D imaging domains, nor different magnification levels as would be present in WSIs. This means that there still exists a lack of large, clinical datasets that are publicly available, which can limit reproducibility and collaboration. Research into directions such as federated learning and differential privacy can be beneficial for solving these challenges [li2025open, wang2024fedmeki].
@@ 6.1.2 Increased Computational Demands of Volumetric Data
For many of the works in this review, it could be observed that larger datasets improved model performance, and that there is a trend towards scaling up. However, scaling is particularly challenging for computationally demanding imaging modalities, such as high-resolution WSIs, and volumetric CT or MRI scans, even outside the medical imaging domain [zuo2024towards]. As foundation models scale up in both data and architecture, ongoing research should prioritize developing more efficient solutions specifically tailored for processing volumetric data, as seen for example in [dominic_improving_2023]
@@ 6.1.3 Pathology FMs Show Plateau in Scaling-Laws
While the challenges in scaling up data and computational resources are still ongoing, pathology FMs have seen many recent successes by scaling up, downstream task performance generally improving with increased data and model size. However, some recent studies show a performance plateau when scaling along these axes [ciga2022self, filiot2023scaling, aben2024towards, campanella2024clinical, chen2024towards, vorontsov2024foundation], which can be attributed to several factors.
@@
First, simply increasing data size may not be enough; better curation is needed to create a more diverse dataset that covers a range of organs, pathologies, and staining types. For example, [dippel2024rudolfv] demonstrated that Virchow’s performance could be matched with a curated dataset containing ten times fewer slides. Furthermore, Atlas [alber2025novel] could even outperform Virchow despite having a much smaller pre-training dataset size, by focusing on including a much broader range of scanners, stainings and magnifications. Similarly, [filiot2024phikon] demonstrated that a smaller ViT-B/16 model pre-trained on 4 million patches using iBOT outperformed Virchow2, a model pre-trained on 350 times more histology images, across three external validation cohorts. This underscores the importance of targeted training strategies over brute-force scaling.
@@
Furthermore, downstream task comparisons often rely on public challenge datasets, but a broader range of benchmark tasks is needed to fully assess pathology FMs. [zimmermann2024virchow] found performance on public tile-level benchmarks plateauing, and [campanella2024clinical] noted more variability in performance when using challenging biomarker prediction tasks as benchmarks. To address these challenges, frameworks such as Patho-Bench [zhang2025accelerating] and eva [gatopoulos2024eva] provide standardized evaluation tools, enabling systematic benchmarking across diverse datasets, tasks, and evaluation protocols.
@@
Finally, current SSL frameworks are optimized for natural images, not pathology images, which have distinct characteristics such as lack of canonical orientation and color variation, as well as interpretation differences depending on the field of view [kang2023benchmarking]. SSL that use for example data augmentations or pretext tasks, such as colorization or rotation prediction frameworks, may need domain-specific adaptations for medical imaging data. [filiot2024phikon] further showed that, while scaling general-purpose FMs like GigaPath, H-Optimus-0, Phikon-v2, UNI, and Virchow2 often benefits downstream tasks, these models do not excel uniformly across all organs or tasks.
Although studies have explored domain-specific adjustments (e.g., weight initialization, data augmentations, magnification levels), both in pathology [wang2022transformer, hua2023pathoduet, kang2023benchmarking, wu2023brow, aben2024towards, ma2024towards, jaume2024transcriptomics, jaume2024multistain, roth2024low, yun2024enhancing, zimmermann2024virchow] and other medical imaging domains [niu_unsupervised_2022, ye_desd_2022, jiang_self-supervised_2022, wu_towards_2023], many open questions remain about developing successful SSL algorithms for medical imaging. As foundation models for all medical imaging domains are increasingly being developed, these questions become relevant for ensuring scalability, generalizability, and the effective utilization of resources across all medical imaging domains.
@@ 6.2.1 Explainability
Foundation models for medical image analysis will need to provide transparent and interpretable results, enabling clinicians and researchers to understand how the model is making predictions and what features are driving those predictions. However, the complexity of foundation models makes it difficult to understand how they make predictions, identify biases, and explain errors, which can have serious consequences in healthcare [marcus2024artificial]. Explainable AI techniques can help address this issue. For example, [abbas_xdecompo_2022] introduced an interpretable decomposition approach for tumor classification, while [pham_i-ai_2023] developed a controllable and interpretable system for chest X-ray diagnoses. [deng_sam-u_2023] add pixel-wise uncertainty estimation to SAM, making it possible to evaluate the the reliability of segmented lesions.
@@ 6.2.2 Robustness and Domain Generalization
Foundation models in medical imaging must remain robust to variations in data acquisition, staining, scanning equipment, and institutional protocols. Without this robustness, models risk producing unreliable or misleading results in real-world deployment. For example, [de2025current] found that many current pathology FMs encode confounding factors such as medical center-specific artifacts more strongly than relevant biological signals. They proposed the robustness index, a quantitative metric to evaluate how well model representations are driven by biological rather than spurious features.
Similarly, [filiot2025distilling] demonstrated that a distilled pathology FM (H0-mini) outperformed larger models on robustness metrics, showing resilience to staining and scanning variability on the PLISM dataset. This highlights that model size does not always correlate with robustness, and that distillation or other targeted training techniques may help improve domain generalization.
@@ 6.2.3 Fairness, Bias and Robustness
Bias presents a significant challenge for FMs, particularly in clinical contexts where the accuracy and fairness of predictions are critical. Models trained on biased datasets risk producing biased predictions, which can disproportionately affect underrepresented populations [glocker_risk_2023]. Addressing this requires diverse and representative training data, alongside robust fairness evaluation frameworks. Studies such as [czum2023bias] and [khan_how_2023] highlight the importance of identifying and mitigating biases in FMs. Specific approaches that aim to reduce bias include cross-lingual pre-training, as proposed by [wan_med-unic_2023], which helps reduce bias in vision-language models. These solutions are vital to ensure FMs provide equitable benefits across diverse patient populations.
Furthermore, ensuring the robustness of foundation models to variables like scanner type, staining protocols, and institutional differences is crucial for their generalizability across clinical settings. [de2025current] found that many current pathology foundation models encode confounding factors, such as medical center-specific staining artifacts, more strongly than relevant biological signals like tissue or cancer type. To address this, they proposed the robustness index, a quantitative metric that measures the extent to which model representations are driven by biological features rather than spurious confounders, providing a practical tool for evaluating model reliability. [filiot2025distilling] explore the distillation of a large foundation model into the much smaller H0-mini. Their model, demonstrates excellent robustness to variations in staining and scanning conditions, significantly outperforming other current pathology FMs.
Finally, with the growing trend of scaling up foundation models, there is a risk of making them inaccessible to institutions with limited computational resources. It should therefore be encouraged to develop more efficient FM architectures to improve accessibility [engelmann2024training, pohjonen2024histoencoderdigitalpathologyfoundation]. Additionally, parameter-efficient fine-tuning strategies, as highlighted by [dutt_parameter-efficient_2023], offer a promising pathway for enabling widespread adoption while minimizing computational demands.
@@ 6.2.4 Regulation
For FMs to be successfully adopted in clinical practice, their risks must be addressed through comprehensive regulation. Medical imaging FMs, like other AI systems, can produce unreliable outputs, such as hallucinations; plausible but incorrect predictions that could lead to misdiagnoses or inappropriate treatments. Recent studies propose methods for detecting hallucinations in medical vision-language FMs [sambara2024radflag, yan2024med], but more research is needed to extend these approaches to vision-based FMs.
Continual learning is a promising research direction to ensure that FMs can adapt to new tasks, disease variants, or data without forgetting previously learned knowledge [wang2024comprehensive, yi2023towards]. This adaptability is crucial in the rapidly evolving medical field.